Find the constants a, b, c if the vector F = (2x+3y+az)i + (bx+2y+3z)j + (2x+cy+3z)k is conservative.


1
Expert's answer
2021-08-28T06:16:34-0400

We require

"\nabla× F=0;"



"0=\begin{vmatrix}
 i & j& k \\
 \frac{\partial}{\partial x}& \frac{\partial}{\partial y}& \frac{\partial}{\partial z}\\
(2x+3y+az)i& (bx+2y+3z)j& (2x+cy+3z)k
\end{vmatrix}"



"0=i(c+1)+j(a-4)+k(b-2);"

"c = -1;\ a = 4;\ b = 2"